# Thinking Outside the (Gray) Box: A Context-Based Score for Assessing Value and Originality in Neural Text Generation

### v2

@@ Introduction
Foundation models (Bommasani et al. 2021), particularly large language models (LLMs) (Bubeck et al. 2023; Gemini Team et al. 2023; Touvron et al. 2023), are significantly transforming creative activities. They can serve as a foundation for co-creation systems involving human and artificial authors (Lin et al. 2023); they can be utilized to generate software code (Rozière et al. 2023); they can even be employed to foster scientific research (Boiko, MacKnight, and Gomes 2023). However, the nature of the self-supervised learning algorithms used for the training of these models tends to make them generate samples as close as possible to the training data distribution (Franceschelli and Musolesi 2024a). In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) (Christiano et al. 2017), is often necessary to generate appropriate and accurate responses. However, this process tends to further reduce output diversity (Kirk et al. 2024), and linguistic creativity tends to be lower than that of humans (Lu et al. 2025). On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request. One typical solution is to sample at a higher temperature to increase diversity. However, this might lead to generating incoherent text (Peeperkorn et al. 2024).
In order to address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs. The definition of CoVO is grounded in the analysis of mutual information (MacKay 2003) between the model’s outputs and inputs, and vice versa. More specifically, we formulate a new optimization problem where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model. In this way, we optimize for solutions that are appropriate for the input request but also different from the outputs we would normally obtain from the model. In particular, we show that our information-theoretic score can be used as a reward in RL-based fine-tuning algorithms, guiding pre-trained models toward more diverse yet valuable solutions. Figure 1 summarizes our proposed approach.
@@
  • We discuss how the score can be practically computed in the case of autoregressive models, and how it can be used as a reward with Group Relative Policy Optimization (GRPO) (Shao et al. 2024), a state-of-the-art reinforcement learning algorithm for fine-tuning LLMs.
@@
  • We evaluate our GRPO-based method on mathematical problem solving, poetry generation, and the tasks included in NoveltyBench (Zhang et al. 2025), demonstrating that our scoring metric correlates with domain-specific measures of value and originality. Moreover, we show that our approach enhances both the quality and diversity of generated outputs, positioning it as a strong candidate for creativity-focused applications of current foundation models.
Related Work
Information Theory and Creativity
@@
The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades. Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli and Musolesi 2024a). However, these methods are often domain-specific and typically require substantial human effort to implement and assess. In contrast, solutions based on information theory (Shannon 1948; Cover 1999) offer a more universally applicable approach.
@@
Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains. Bayesian surprise (Baldi and Itti 2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure either novelty (França et al. 2016; Varshney et al. 2019) or surprise (Mazzaglia et al. 2022; Schmidhuber 2010). Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures. Surprisal (Tribus 1961), i.e., Shannon’s self-information, has also been used (Bunescu and Uduehi 2019; Fernandez Monsalve, Frank, and Vigliocco 2012); Barto, Mirolli, and Baldassarre (2013) extensively discuss surprisal and Bayesian surprise, and how novelty differs from them. Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more (Achiam and Sastry 2016). Sun et al. (2025) apply this idea to improve exploration in RLHF (Christiano et al. 2017). Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience. Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness (Li et al. 2016). However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
LLMs and Creativity
@@
Since the introduction of GPT models (Brown et al. 2020; OpenAI 2023) and their competitors (e.g., Touvron et al. 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this (Franceschelli and Musolesi 2024b). For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs (Stevenson et al. 2022) and to investigate methods for enhancing their performance (Goes et al. 2023; Summers-Stay, Voss, and Lukin 2023). Porter and Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones. In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal. Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems (Popescu-Belis et al. 2023) or utilized in zero-shot settings to emulate the writing styles of famous authors (Sawicki et al. 2023). It has also been shown that these models can be fine-tuned via RLHF (Christiano et al. 2017) to write short poems that human evaluators find more creative (Pardinas et al. 2023). Reinforcement learning (RL) can also encourage large language models (LLMs) to produce more diverse outputs (Chung et al. 2025), and can be leveraged to optimize metrics commonly linked to creative expression (Ismayilzada et al. 2025). Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human (Ding et al. 2023) or AI (Bradley et al. 2024) feedback to measure the quality of the generated outputs.
Preliminaries
Language Modeling
@@
A $\boldsymbol{\theta}$-parameterized autoregressive language model is a probability distribution $p_{\boldsymbol{\theta}}(\mathbf{x})$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$ of size $N$. The probability distribution is factorized as $p_{\boldsymbol{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}})$, where $\mathbf{x_{<t}}=x_{1}\ldots x_{t-1}$. The language model is usually trained to maximize the likelihood of the true distribution $p^{*}(\mathbf{x})$ for any $\mathbf{x}$ from a reference dataset (the training set). In other words, given an input $\mathbf{x_{<t}}$, the model learns to approximate the probability of each token from $\mathcal{V}$ being $x_{t}$. While this makes such a model immediately capable of scoring the probability of a given text, it also allows for the generation of new sentences. Given a conditional input (prompt) $\mathbf{z}=(z_{1}\ldots z_{L})$, we can decode $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})$ as the continuation of $\mathbf{z}$, i.e., through the factorized representation $p_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}p_{\boldsymbol{\theta}}(x_{t}|\mathbf{x_{<t}},\mathbf{z})$.
Reinforcement Learning for Language Models
@@
Due to its adherence to the formal framework of Markov decision processes (Sutton and Barto 2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation (Bachman and Precup 2015). The LLM plays the role of the agent, and each generated token represents an action $a_{t}$. The current version of the generated output $\mathbf{x_{t}}$ is part of the state $\mathbf{s_{t}}$ (potentially with additional information such as initial prompts). Finally, the reward $r_{t+1}$ measures the “quality” of the current output. A common strategy is to assign a zero reward for each $\mathbf{x}_{t},t\neq T$ and a sentence-based reward when the final output is generated. Within this framework, any policy-based method can be employed to train or fine-tune the LLM to optimize a given objective. Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences (Franceschelli and Musolesi 2024c).
@@
A widely used RL algorithm for fine-tuning LLMs is Proximal Policy Optimization (PPO) (Schulman et al. 2017), which aims to maximize the following objective:
@@
$\mathcal{J}(\boldsymbol{\theta})=\mathbb{E}_{t}\!\left[\min(r_{t}(\boldsymbol{\theta})\hat{A}_{t},\mathtt{clip}(r_{t}(\boldsymbol{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$ (1)
@@
where $r_{t}(\boldsymbol{\theta})=\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}{\pi_{\boldsymbol{\theta}_{old}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}$ with $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}_{old}}$ denoting the current and old policy models, repectively; $\mathbf{x}$ is the output sampled from the old policy given the prompt $\mathbf{z}$; and $\epsilon$ is a clipping factor used to stabilize training. The advantage $\hat{A}_{t}$ is usually computed through Generalized Advantage Estimation (Schulman et al. 2016) based on the full rewards $R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{\pi_{\boldsymbol{\theta}}(\mathbf{x}|\mathbf{z})}{\pi_{ref}(\mathbf{x}|\mathbf{z})}$, thus integrating a KL penalty with respect to a reference model (usually, the same model before fine-tuning), and a learned value function $v_{\boldsymbol{\phi}}(\mathbf{s})$. However, learning such a value function is computationally intensive, and its training is complicated by the fact that only the last state is scored by the reward function. Moreover, the inclusion of the KL penalty as an auxiliary reward term adds complexity to the advantage estimation process. To address these issues, Group Relative Policy Optimization (GRPO) (Shao et al. 2024) has been introduced. GRPO directly adds a KL divergence term $-\beta D_{KL}(\pi_{\boldsymbol{\theta}}||\pi_{ref})$ to the loss rather than to the single rewards, and especially obviates the need for a value function approximator by using the average reward of multiple sampled outputs as the baseline:
@@
A Context-Based Score for Valuable and Original Generation
@@
$I(x,y)=h(x)-h(x|y)=h(y)-h(y|x)$ (3)
where the entropy is $h(a)=-\log p(a)$, therefore:
$I(x,y)$ $=\log p(x|y)-\log p(x)$ (4)
$=\log p(y|x)-\log p(y).$
@@
$I(\mathbf{x},\mathbf{y},\lambda_{1},\lambda_{2})$ $=$ $\lambda_{1}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{y}|\mathbf{x})-$ (7)
$\lambda_{2}\log p(\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})$
$=$ $(\lambda_{1}-\lambda_{2})\log p(\mathbf{y}|\mathbf{x})+$
$\lambda_{2}\log p(\mathbf{x}|\mathbf{y})-\lambda_{2}\log p(\mathbf{x}).$
@@
Let us now consider the case where $\lambda_{v},\lambda_{o}>0$, for example, $\lambda_{v}=\lambda_{o}=1$. Solving this maximization problem involves finding the target $\mathbf{y}$ that maximizes the posterior probability of $\mathbf{x}$ while also being unlikely given $\mathbf{x}$. In other words, the optimal $\mathbf{y}^{*}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$. While $-\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal (Tribus 1961), is widely used to measure diversity and surprise (Barto, Mirolli, and Baldassarre 2013), the other term, $\log p(\mathbf{x}|\mathbf{y})$, can be used to measure value and effectiveness. Indeed, if the request (e.g., a problem or a task) can be easily predicted from the outcome, the outcome must be a (good) example of that task or a correct solution for that problem. While other prominent definitions, such as Boden’s, seek to include a third requirement for creativity, i.e., novelty (Boden 2003), we believe that by considering only a specific context, novelty and surprise become indistinguishable. In particular, being novel in a specific context $\mathbf{x}$ means doing something different from what has been previously experienced under $\mathbf{x}$ by the creator. However, since a self-supervised-trained model can be seen as a compression of training data (Franceschelli, Cevenini, and Musolesi 2024), $-\log p(\mathbf{y}|\mathbf{x})$ becomes not only a measure of unexpectedness but also of novelty. Therefore, we choose to refer to this term as originality, as it encompasses both novelty and surprise. On the other hand, $\log p(\mathbf{x}|\mathbf{y})$ is less ambiguous and maps directly into value.
@@
Implementation and Optimization with Autoregressive Models

### v3

@@ 1 Introduction
Foundation models (Bommasani et al., 2021), particularly large language models (LLMs) (Gemini Team et al., 2023; Guo et al., 2025; Touvron et al., 2023), are significantly transforming creative activities. They can serve as a foundation for co-creation systems involving human and artificial authors (Lin et al., 2023), can be utilized to generate software code (Rozière et al., 2023), or even to foster scientific research (Boiko et al., 2023). However, the nature of the self-supervised learning algorithms used for the training of these models tends to make their sampling distribution as close as possible to the training data distribution (Franceschelli & Musolesi, 2024a). In addition, fine-tuning, such as that based on reinforcement learning from human feedback (RLHF) (Christiano et al., 2017), is often necessary to generate appropriate and accurate responses. However, this process tends to reduce output diversity further (Kirk et al., 2024), and linguistic creativity tends to be lower than that of humans (Lu et al., 2025). On the contrary, LLMs for creative tasks should produce more novel and surprising texts that maintain a high level of correctness and adherence to the request. One typical solution is to sample at a higher temperature to increase diversity. However, this might lead to generating less coherent text (Peeperkorn et al., 2024).
To address the issues described above, we propose a new training approach for creative tasks based on CoVO, a Context-based score for Value and Originality, with the goal of taking into consideration both value and originality of the neurally-generated text in the optimization of LLMs. The definition of CoVO is grounded in the analysis of mutual information (MacKay, 2003) between the model’s outputs and inputs, and vice versa. More specifically, we formulate a new optimization objective where, given a specific input, the desired output is derived by simultaneously maximizing the conditional probability of the input given the output and minimizing the conditional probability of the output given the input under the generative model. In this way, we optimize for solutions that are appropriate for the input request but also different from the outputs we would normally obtain from the model. In particular, we show that our information-theoretic score can be used as a reward in RL-based fine-tuning algorithms, guiding pre-trained models toward more diverse yet valuable solutions. Figure 1 summarizes our proposed approach.
@@
  • We discuss how the score can be practically computed in the case of autoregressive models, and how it can be used as a reward with Group Relative Policy Optimization (GRPO) (Shao et al., 2024), a state-of-the-art reinforcement learning algorithm for fine-tuning LLMs.
@@
  • We evaluate our GRPO-based method on mathematical problem solving, poetry generation, and the tasks included in NoveltyBench (Zhang et al., 2025), demonstrating that our approach can enhance both the quality and diversity of generated outputs, positioning it as a strong candidate for creativity-focused applications of current foundation models.
2 Related Work
2.1 Information Theory and Creativity
@@
The quest to provide a mathematical and computational definition of creativity has been a significant focus in recent decades. Numerous methods have been developed to define various dimensions or attributes for evaluating the creativity of AI-generated products (see, for example, Franceschelli & Musolesi, 2024a). However, these methods are often domain-specific and typically require substantial human effort to implement and assess. In contrast, solutions based on information theory (Shannon, 1948; Cover, 1999) offer a more universally applicable approach.
@@
Information-theoretic methods can quantify creativity by measuring the novelty and complexity of generated outputs, without the need for extensive human intervention, making them suitable for a wide range of domains. Bayesian surprise (Baldi & Itti, 2010), i.e., the divergence between a prior and a posterior belief, has been extensively used to measure different shades of originality, such as novelty (França et al., 2016; Varshney et al., 2019) and surprise (Mazzaglia et al., 2022; Schmidhuber, 2010). Nevertheless, Varshney (2019) demonstrated that there is a mathematical limit for Bayesian surprise when combined with quality measures. Surprisal (Tribus, 1961), i.e., Shannon’s self-information, has also been used (Bunescu & Uduehi, 2019; Fernandez Monsalve et al., 2012); Barto et al. (2013) extensively discuss surprisal, Bayesian surprise, and novelty. Crucially, in the context of RL, surprisal has been used as a form of intrinsic motivation to encourage the agent to explore more (Achiam & Sastry, 2016). Sun et al. (2025) apply this idea to improve exploration in RLHF (Christiano et al., 2017). A similar strategy can be applied during LLM fine-tuning, either by explicitly maximizing the model’s perplexity (Dai et al., 2025) or by decoupling entropy and cross-entropy from KL regularization and assigning greater weight to the former (Slocum et al., 2025). Burns (2006) proposes to use entropy for expectation and violation, plus posterior probability for explanation in the context of aesthetic experience. Additionally, mutual information has been applied to neural conversation models to improve both diversity and appropriateness (Li et al., 2016). However, all these existing approaches are not able to capture and simultaneously optimize value and originality at the same time.
2.2 LLMs and Creativity
@@
Since the introduction of GPT models (Brown et al., 2020; OpenAI, 2023) and their competitors (e.g., Touvron et al., 2023), researchers have been keenly exploring the potential for LLMs to exhibit creativity and the methods to achieve this (Franceschelli & Musolesi, 2025). For example, human creativity tests like the Alternate Uses Test have been employed to evaluate the creativity of LLMs (Stevenson et al., 2022) and to investigate methods for enhancing their performance (Goes et al., 2023; Summers-Stay et al., 2023). Porter & Machery (2024) report that non-expert poetry readers already favor AI-generated poems over human-authored ones. In contrast, Davis (2024) argues that ChatGPT’s poetry is incompetent and banal. Either way, instead of being used off-the-shelf, LLMs can be fine-tuned to produce more rhyming poems (Popescu-Belis et al., 2023) or utilized in zero-shot settings to emulate the writing styles of famous authors (Sawicki et al., 2023). It has also been shown that these models can be fine-tuned via RLHF (Christiano et al., 2017) to write short poems that human evaluators find more creative (Pardinas et al., 2023). Reinforcement learning (RL) can also encourage large language models (LLMs) to produce more diverse outputs (Chung et al., 2025), and can be leveraged to optimize metrics commonly linked to creative expression (Ismayilzada et al., 2025). Finally, it is possible to leverage quality-diversity algorithms to generate more creative products; these methods can be based on human (Ding et al., 2023) or AI (Bradley et al., 2024) feedback to measure the quality of the generated outputs.
3 Preliminaries
3.1 Language Modeling
@@
A $\bm{\theta}$-parameterized autoregressive language model is a probability distribution $p_{\bm{\theta}}(\mathbf{x})$ over a variable-length text sequence $\mathbf{x}=(x_{1}\ldots x_{T})$, where $T$ is the sequence length and each token $x_{t}$ is in a finite vocabulary $\mathcal{V}$ of size $N$. The probability distribution is factorized as $p_{\bm{\theta}}(\mathbf{x})=\prod_{t=1}^{T}p_{\bm{\theta}}(x_{t}|\mathbf{x_{<t}})$, where $\mathbf{x_{<t}}=x_{1}\ldots x_{t-1}$. The language model is usually trained to maximize the likelihood of the true distribution $p^{*}(\mathbf{x})$ for any $\mathbf{x}$ from a reference dataset (the training set). In other words, given an input $\mathbf{x_{<t}}$, the model learns to approximate the probability of each token from $\mathcal{V}$ being $x_{t}$. While this makes such a model immediately capable of scoring the probability of a given text, it also allows for the generation of new sentences. Given a conditional input (prompt) $\mathbf{z}=(z_{1}\ldots z_{L})$, we can decode $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})$ as the continuation of $\mathbf{z}$, i.e., through the factorized representation $p_{\bm{\theta}}(\mathbf{x}|\mathbf{z})=\prod_{t=1}^{T}p_{\bm{\theta}}(x_{t}|\mathbf{x_{<t}},\mathbf{z})$.
3.2 Reinforcement Learning for Language Models
@@
Due to its adherence to the formal framework of Markov decision processes (Sutton & Barto, 2018), RL can be used as a solution to the generative modeling problem in the case of autoregressive tasks such as text generation (Bachman & Precup, 2015). The LLM plays the role of the agent, and each generated token represents an action $a_{t}$. The current version of the generated output $\mathbf{x_{t}}$ is part of the state $\mathbf{s_{t}}$ (potentially with additional information such as initial prompts). Finally, the reward $r_{t+1}$ measures the “quality” of the current output. A common strategy is to assign a zero reward for each $\mathbf{x}_{t},t\neq T$ and a sentence-based reward when the final output is generated. Within this framework, any policy-based method can be employed to train or fine-tune the LLM to optimize a given objective. Indeed, RL facilitates the use of non-differentiable reward functions, enabling the optimization of test-time metrics, domain-specific targets, and human preferences (Franceschelli & Musolesi, 2024b).
@@
A widely used RL algorithm for fine-tuning LLMs is Proximal Policy Optimization (PPO) (Schulman et al., 2017), which aims to maximize the following objective:
@@
$\mathcal{J}(\bm{\theta})=\hat{\mathbb{E}}_{t}\!\left[\min(r_{t}(\bm{\theta})\hat{A}_{t},\mathtt{clip}(r_{t}(\bm{\theta}),1-\epsilon,1+\epsilon)\hat{A}_{t})\right]$ (1)
@@
where $r_{t}(\bm{\theta})=\frac{\pi_{\bm{\theta}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}{\pi_{\bm{\theta}_{old}}(\mathbf{x}_{t}|\mathbf{z},\mathbf{x}_{<t})}$ with $\pi_{\bm{\theta}}$ and $\pi_{\bm{\theta}_{old}}$ denoting the current and old policy models, repectively; $\mathbf{x}$ is the output sampled from the old policy given the prompt $\mathbf{z}$; and $\epsilon$ is a clipping factor used to stabilize training. The advantage $\hat{A}_{t}$ is usually computed through Generalized Advantage Estimation (Schulman et al., 2016) based on the full rewards $R(\mathbf{z},\mathbf{x})=r(\mathbf{z},\mathbf{x})-\beta\log\frac{\pi_{\bm{\theta}}(\mathbf{x}|\mathbf{z})}{\pi_{ref}(\mathbf{x}|\mathbf{z})}$, thus integrating a KL penalty with respect to a reference model (usually, the same model before fine-tuning), and a learned value function $v_{\bm{\phi}}(\mathbf{s})$. However, learning such a value function is computationally intensive, and its training is complicated by the fact that only the last state is scored by the reward function. Moreover, the inclusion of the KL penalty as an auxiliary reward term adds complexity to the advantage estimation process. To address these issues, Group Relative Policy Optimization (GRPO) (Shao et al., 2024) has been introduced. GRPO directly adds a KL divergence term $\beta D_{KL}(\pi_{\bm{\theta}}||\pi_{ref})$ to the loss rather than to the single rewards, and especially obviates the need for a value function approximator by using the average reward of multiple sampled outputs as the baseline:
@@
4 A Context-Based Score for Valuable and Original Generation
@@
I(x,y)=h(x)h(x|y)=h(y)h(y|x)I(x,y)=h(x)-h(x|y)=h(y)-h(y|x) (3)
where the self-information is $h(a)=-\log p(a)$, therefore:
$I(x,y)=\log p(x|y)-\log p(x)=\log p(y|x)-\log p(y).$ (4)
@@
$I(\mathbf{x},\mathbf{y},\lambda_{1},\lambda_{2})$ $=$ $\lambda_{1}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{y}|\mathbf{x})-\lambda_{2}\log p(\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})$ (7)
$=$ $(\lambda_{1}-\lambda_{2})\log p(\mathbf{y}|\mathbf{x})+\lambda_{2}\log p(\mathbf{x}|\mathbf{y})-\lambda_{2}\log p(\mathbf{x}).$
@@
Let us now consider the case where $\lambda_{v},\lambda_{o}>0$, for example, $\lambda_{v}=\lambda_{o}=1$. Solving this maximization problem involves finding the target $\mathbf{y}$ that maximizes the posterior probability of $\mathbf{x}$ while also being unlikely given $\mathbf{x}$. In other words, the optimal $\mathbf{y}^{*}$ must be unexpected and diverse from $p(\mathbf{y}|\mathbf{x})$, but it must also be explainable by $\mathbf{x}$. $\log p(\mathbf{y}|\mathbf{x})$, commonly known as surprisal (Tribus, 1961), is widely used to measure diversity and surprise (Barto et al., 2013), and adheres to the first requirement from the standard definition of creativity by Runco & Jaeger (2012), i.e., originality. Conversely, $\log p(\mathbf{x}|\mathbf{y})$ can be used to measure value or effectiveness, the second requirement of the definition. If the request (e.g., a problem or task) can be inferred from the outcome, the latter constitutes an appropriate instance of that task or a correct, useful solution to that problem (e.g., if the request is for a sonnet or a sci-fi screenplay, the generated artifact is valuable if identified as a poem satisfying the metrical constraints of a sonnet or as a screenplay adhering to a sci-fi theme).
@@
5 Implementation and Optimization with Autoregressive Models